Consider a money demand function that takes the form (M/P)^d = Y/3i, where M is the quantity of money, P is the price lelve, Y is real output, and i is the nominal interest rate (measured in percentage points.)

a) If the economy is growing g% per year and velocity is growing twice as fast what are the growth rates in real money balances and the interest rate?

b) Supposing the money supply is growing at 3% per year, what values of g indicate inflation? Deflation?

I am struggling with these questions. Could you please answer the following questions. Thank you
Answer #1

The Velocity of Money Velocity of money: the rate at which money changes hands - Notation: - nominal GDP - (price level) x (real GDP) M -money supply V velocity Velocity formula: V- CHAPTER 22 MONEY GROWTH AND INFLATION 20

Real money demand:

small (rac{M}{P})^d = rac{Y}{3i}small = rac{Y}{v} where v = 3i = velocity of money

small ln(rac{M}{P})^d = lnrac{Y}{v} = ln Y - ln v

Differentiating both sides:

small Rightarrow rac{(Delta rac{M}{P})^d}{( rac{M}{P})^d} = rac{Delta Y}{Y} - rac{Delta v}{v}

If the economy is growing g% per year and velocity is growing twice as fast:

small rac{Delta Y}{Y} = g% and = 200%

small Rightarrow rac{(Delta rac{M}{P})^d}{( rac{M}{P})^d} = rac{Delta Y}{Y} - rac{Delta v}{v}(9-200)%

Also

small 3i = v Rightarrow i = rac{v}{3}

Thus,

200/3-66.67% = = 3

b)

If money supply grows at 3% per year,

small P = MV/Y

small ln P = ln (MV/Y) = ln M + ln V - ln Y

Differentiating both sides we get:

small rac{Delta P}{P} = rac{Delta M}{M}+rac{Delta V}{V}-rac{Delta Y}{Y}

small Rightarrow rac{Delta P}{P} = (3+203-g)% = (203-g)%

For inflation, small rac{Delta P}{P} > 0 Rightarrow (203-g) > 0 Rightarrow g < 203

For deflation, small rac{Delta P}{P} < 0 Rightarrow (203-g) < 0 Rightarrow g > 203